OpenStudy (anonymous):

Find a fourth degree polynomial in Z2[x] whose roots are the four elements of the field Z2[x]/(x^2+x+1) (given to be [0] [1] [x] [x+1])

OpenStudy (anonymous):

note that by Z2[x] I mean 

OpenStudy (anonymous):

i think this might be easier than it looks

OpenStudy (anonymous):

Well I think this has to do with something called an extension field, which the textbook did not go over before this question appeared (its in the next section). Also I do not seem to understand it cause I keep getting what appears to be an incorrect answer

OpenStudy (anonymous):

if i am reading it correctly, you only have 3 elements of \(\mathbb{Z}_2[x]/x^2+x+1\) namely \(\{0,a,1,a+1\}\) the equivalence classes or in the square bracket notation \[\{[0],[1],[x],[x+1]\}\]

OpenStudy (anonymous):

make that 3 non zero elements!

OpenStudy (anonymous):

these elements form a group under multiplication, and since the order of the group is 3, then each has order 3 that means in my notation \(a^3=1\) for all or \(a^3-1=0\) so your polynomial is \(x^3-1\) in order to add the zero element, the polynomial will be \(x^4-x\)

OpenStudy (anonymous):

so can you take me through the process of getting 0 by plugging in [x] then? I keep getting that it is [x+1] which is not 0

OpenStudy (anonymous):

the equivalence class that i wrote as \(a\) and you have written as \([x]\) is the root of that polynomial \(x^2+x+1\)

OpenStudy (anonymous):

the other root is \([x+1]\) and that is an easy check, because \[f(a+1))=(a+1)^2+a+1+1=a^2+2a+1+a+1+1=a^2+a+1=0\]

OpenStudy (anonymous):

ooops that kind of ran over \[f(a+1)=a^2+2a+1+a+1+1\] \[=a^2+a+1=0\] because \(a\) is a root of \(x^2+x+1\) and \(2, 2a=0\)

OpenStudy (anonymous):

yes but when I plug it into x^4 + x in I just get [x^4 + x} which (if I did my division right...is equivalent to [x+1]

OpenStudy (anonymous):

I must be doing my division wrong trying to get the equivalent class....

OpenStudy (anonymous):

maybe not exactly sure

OpenStudy (anonymous):

did you do an example in class where you take \[\mathbb{R}/x^2+1\]?

OpenStudy (anonymous):

you are introducing other zeros in to \(\mathbb{R}\) namely \(i\) and \(-i\) this is the same kind of idea i think

OpenStudy (anonymous):

Sounds familar, alright thank you

OpenStudy (anonymous):

yw